OpenStudy (anonymous):

Calculate the length of the third side of the triangle; 4cm and 3 cm

OpenStudy (primeralph):

if it's a right triangle, answer is 5.

OpenStudy (anonymous):

Thanks, it is not drawn to scale! How did u get 5?

OpenStudy (anonymous):

a^2 + b^2 = c^2 where c is the longest side this special case of cosine law only works for RIGHT ANGLE triangles also, you should remember that a 3,4,5 triangle exists :)

OpenStudy (anonymous):

How would I work out 12cm by 9cm

OpenStudy (espex):

Assuming the triangle in your second query is a right triangle, you would use the same Pythagorean theorem of \(a^2+b^2=c^2\) and solve for the missing side. 

OpenStudy (anonymous):

Just to make sure that the side lengths that you came up with for a certain triangle are in-fact correct, check using the Triangle In-equality:Assume triangle abc to be any generic triangle. By the Triangle Inequality, we have:\[\bf a + b > c, \ b+c > a, \ a + c > b\], where a, b, c are the sides of the triangle. This just means that any two sides of a triangle will always be greater than the third one. If you ever feel that the lengths you have are incorrect, just check to see if the sum of two sides is greater than the third one. If so, then your side lengths should work. @deemills